OpenStudy (anonymous):

A rock falls from rest a vertical distance of 0.57 meters to the surface of a planet in 0.59 seconds. What is the magnitude of the acceleration due to gravity on that planet? [Answer to 3 decimal places. No units]

OpenStudy (anonymous):

x=1/2at^2

OpenStudy (anonymous):

.57=1/2a(.59)^2 solve for a which give you acceleration due to gravity

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

i thought i could divide the distance by time to get velocity and divide that by time again to get acceleration.

OpenStudy (anonymous):

Thanks for the help.

OpenStudy (anonymous):

h = ut + 1/2 gt^2 here, h=0.57, u=o,t=0.59, g=? substituting, we get g = 3.27